# A convex formulation of covariate-adjusted Gaussian graphical models via natural parametrization

### v1

@@ 5 Simulation Study
For each simulation setting, we generate a sparse $\boldsymbol{\mathbf{\Gamma}}\in\mathbb{R}^{p\times q}$ by independently determining if $\Gamma_{jk}$ is non-zero with probability $0.3$. The non-zero entries are sampled from $N(0,1)$. Then $\boldsymbol{\mathbf{\Gamma}}$ is scaled by a constant to ensure that the signal-to-noise ratio is approximately $1$ for each observation.
We generate the population graph $\boldsymbol{\mathbf{B}}_{0}$ using a preferential attachment algorithm [Barabási and Albert, 1999] with power $1$. We select $q_{e}=5$ out of $q$ covariates to have nonzero effects and generate the covariate-adjusted components $\boldsymbol{\mathbf{B}}_{h}$ for $h\in[q_{e}]$ as Erdős-Renyi graphs [Erdős and Rényi, 1959] with edge probability $v_{e}=0.01$. This choice of graph structure ensures that the covariate-dependent graphs are more sparse than the population graph; see Clauset et al. [2009] in particular for a discussion of power law graph models in applications. After determining the graph structure, we generate entries in $\{\boldsymbol{\mathbf{B}}_{h}\}$ by independently and uniformly sampling from $[-0.5,-0.35]\cup[0.35,0.5]$ and then dividing each row $j$ by $\sum_{h}\sum_{k\neq j}\lvert\beta_{jkh}\rvert\times 1.5$ before symmetrizing each $\{\boldsymbol{\mathbf{B}}_{h}\}$ by taking the average of $\beta_{jkh}$ and $\beta_{kjh}$. These steps help ensure a diagonally dominant precision matrix, and mirror the data model of Zhang and Li [2022].
We select half of the $q$ covariates to be binary and half to be continuous from $\operatorname{Unif}(0,1)$ and then standardize the continuous covariates across the $n$ observations to have zero mean and unit variance. With $\boldsymbol{\mathbf{\Gamma}}$ and $\{\boldsymbol{\mathbf{B}}_{h}\}$ in hand, we generate a precision matrix $\boldsymbol{\mathbf{\Omega}}^{(i)}=\boldsymbol{\mathbf{B}}_{0}+\sum_{h=1}^{q}\boldsymbol{\mathbf{B}}_{h}u_{h}^{(i)}$. Then we generate a dataset under the original model (3) by setting $\boldsymbol{\mathbf{\mu}}^{(i)}=\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}}^{(i)}$ and and a second dataset under (7), termed the “natural model”, via $\boldsymbol{\mathbf{\mu}}^{(i)}=\boldsymbol{\mathbf{\Sigma}}^{(i)}\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}}^{(i)}$. Finally, we sample a vector of responses $\boldsymbol{\mathbf{x}}^{(i)}\sim N_{p}(\boldsymbol{\mathbf{\mu}}^{(i)},\boldsymbol{\mathbf{\Sigma}}^{(i)})$ for $i\in[n]$. In both models, we set $\operatorname{diag}(\boldsymbol{\mathbf{\Omega}}^{(i)})=1$ so that $\sigma^{2}_{\varepsilon_{j}}=1$.
For each of the settings $p=25,q=50$ and $p=25,q=100$, we generate $100$ independent data sets of $n=200$ and $n=400$ samples and run both RegGMM and cspine on each data set. For a fair comparison, both methods have two tuning parameters selected via $5$-fold cross-validation on a path of $100$ $\lambda$ parameters and $10$ mixture parameters. With the notation in (16), this means cross-validating over $100$ values of $\lambda_{0}$ and $\alpha_{s}=0.1,0.2,\dotsc,1$.
We report the mean and standard error of the following metrics after applying (13): TPR, the true positive rate of detected edges across all $\{\boldsymbol{\mathbf{B}}_{h}\}$; $\text{TPR}_{\text{pop}}$, the true positive rate of detected edges in the population graph $\boldsymbol{\mathbf{B}}_{0}$; $\text{FPR}_{\text{pop}}$, the false positive rate of detected edges in $\boldsymbol{\mathbf{B}}_{0}$; $\text{TPR}_{\text{cov}}$, the true positive rate of detected edges in covariate graphs $\boldsymbol{\mathbf{B}}_{h}$, $h\in[q]$; the nodewise estimation error given by $\boldsymbol{\mathbf{\beta}}_{\text{err}}=\sum_{j=1}^{p}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{j}-\boldsymbol{\mathbf{\beta}}_{j}\rVert_{2}$; and the average error of the estimated precision matrix given by $\boldsymbol{\mathbf{\Omega}}_{\text{err}}=\sum_{i=1}^{n}\lVert{\hat{\boldsymbol{\mathbf{\Omega}}}}^{(i)}-\boldsymbol{\mathbf{\Omega}}^{(i)}\rVert^{2}_{F,\text{off}}/n$, where the norm is taken over off-diagonal entries. The overall false positive rate of detected edges is less than $0.005$ in all cases.
@@
In Table 1, we see that cspine outperforms RegGMM under the natural model setting across all metrics. Table 2 compares the two methods under the setting (3) where RegGMM is correctly specified and cspine is misspecified. We see that despite the misspecification, cspine is competitive with RegGMM in edge detection and estimation error due to the convex formulation of the nodewise regressions. Both methods are conservative in their detection of edges in the covariate networks.
$n$ $(p,q)$ Method TPR $\text{TPR}_{\text{pop}}$ $\text{FPR}_{\text{pop}}$ $\text{TPR}_{\text{cov}}$ $\boldsymbol{\mathbf{\beta}}_{\text{err}}$ $\boldsymbol{\mathbf{\Omega}}_{\text{err}}$
200 (25, 50) cspine $\mathbf{0.73}$ $(0.08)$ $\mathbf{0.89}$ $(0.07)$ $\mathbf{0.07}$ $(0.02)$ $\mathbf{0.45}$ $(0.21)$ $\mathbf{5.66}$ $(0.50)$ $\mathbf{1.09}$ $(0.15)$
RegGMM $0.62$ $(0.09)$ $0.80$ $(0.10)$ $0.16$ $(0.06)$ $0.30$ $(0.18)$ $7.43$ $(0.61)$ $1.82$ $(0.26)$
(25, 100) cspine $\mathbf{0.66}$ $(0.10)$ $\mathbf{0.82}$ $(0.09)$ $\mathbf{0.05}$ $(0.02)$ $\mathbf{0.40}$ $(0.24)$ $\mathbf{5.98}$ $(0.55)$ $\mathbf{1.22}$ $(0.17)$
RegGMM $0.55$ $(0.08)$ $0.71$ $(0.08)$ $0.14$ $(0.05)$ $0.28$ $(0.19)$ $7.73$ $(0.56)$ $1.95$ $(0.28)$
400 (25, 50) cspine $\mathbf{0.77}$ $(0.10)$ $\mathbf{0.96}$ $(0.04)$ $\mathbf{0.09}$ $(0.02)$ $\mathbf{0.52}$ $(0.24)$ $\mathbf{4.50}$ $(0.59)$ $\mathbf{0.66}$ $(0.08)$
RegGMM $0.67$ $(0.08)$ $0.86$ $(0.07)$ $0.17$ $(0.07)$ $0.41$ $(0.20)$ $5.99$ $(0.55)$ $1.22$ $(0.17)$
(25, 100) cspine $\mathbf{0.72}$ $(0.09)$ $\mathbf{0.94}$ $(0.05)$ $\mathbf{0.08}$ $(0.02)$ $\mathbf{0.44}$ $(0.21)$ $\mathbf{4.96}$ $(0.60)$ $\mathbf{0.76}$ $(0.09)$
RegGMM $0.62$ $(0.08)$ $0.85$ $(0.08)$ $0.22$ $(0.07)$ $0.32$ $(0.17)$ $6.80$ $(0.58)$ $1.40$ $(0.20)$
@@
Table 1: Mean and standard error of performance metrics over $100$ data sets.
$n$ $(p,q)$ Method TPR $\text{TPR}_{\text{pop}}$ $\text{FPR}_{\text{pop}}$ $\text{TPR}_{\text{cov}}$ $\boldsymbol{\mathbf{\beta}}_{\text{err}}$ $\boldsymbol{\mathbf{\Omega}}_{\text{err}}$
200 (25, 50) cspine $\mathbf{0.53}$ $(0.06)$ $\mathbf{0.77}$ $(0.08)$ $\mathbf{0.12}$ $(0.02)$ $0.09$ $(0.10)$ $\mathbf{7.70}$ $(0.29)$ $\mathbf{2.18}$ $(0.24)$
RegGMM $0.50$ $(0.07)$ $0.70$ $(0.10)$ $0.14$ $(0.05)$ $\mathbf{0.12}$ $(0.11)$ $8.02$ $(0.41)$ $2.20$ $(0.25)$
(25, 100) cspine $\mathbf{0.44}$ $(0.07)$ $\mathbf{0.67}$ $(0.09)$ $\mathbf{0.09}$ $(0.02)$ $0.04$ $(0.06)$ $\mathbf{7.97}$ $(0.32)$ $\mathbf{2.43}$ $(0.25)$
RegGMM $0.39$ $(0.07)$ $0.59$ $(0.12)$ $0.19$ $(0.07)$ $\mathbf{0.05}$ $(0.06)$ $8.80$ $(0.45)$ $2.69$ $(0.32)$
400 (25, 50) cspine $\mathbf{0.60}$ $(0.05)$ $\mathbf{0.92}$ $(0.05)$ $0.18$ $(0.03)$ $0.12$ $(0.10)$ $6.71$ $(0.28)$ $1.54$ $(0.18)$
RegGMM $0.59$ $(0.07)$ $0.80$ $(0.10)$ $\mathbf{0.14}$ $(0.05)$ $\mathbf{0.28}$ $(0.15)$ $\mathbf{6.64}$ $(0.42)$ $\mathbf{1.49}$ $(0.19)$
(25, 100) cspine $\mathbf{0.55}$ $(0.06)$ $\mathbf{0.88}$ $(0.07)$ $\mathbf{0.15}$ $(0.02)$ $0.12$ $(0.10)$ $\mathbf{6.83}$ $(0.26)$ $\mathbf{1.69}$ $(0.22)$
RegGMM $\mathbf{0.55}$ $(0.07)$ $0.81$ $(0.09)$ $0.22$ $(0.07)$ $\mathbf{0.19}$ $(0.14)$ $7.15$ $(0.42)$ $\mathbf{1.69}$ $(0.23)$
@@
Table 2: Performance metrics over $100$ data sets under specification (3).
Next, we explore the error bound in Theorem 1 through simulations. We fixed $p=25$, $q=10$, and $n=300$. In the first set of simulations, we fixed a collection $\{\boldsymbol{\mathbf{B}}_{h}\}$ and generated $300$ data sets from $\boldsymbol{\mathbf{\Gamma}}$ matrices of varying sparsities. In the second set of simulations, we fixed a sparse $\boldsymbol{\mathbf{\Gamma}}$ and generated data sets from $300$ different collections of $\{\boldsymbol{\mathbf{B}}_{h}\}$ with varying sparsity levels. In Figure 2, we plot the error $\boldsymbol{\mathbf{\gamma}}_{\text{err}}+\boldsymbol{\mathbf{\beta}}_{\text{err}}$ where $\boldsymbol{\mathbf{\gamma}}_{\text{err}}=\sum_{j=1}^{p}\lVert\hat{\boldsymbol{\mathbf{\gamma}}}_{j}-\boldsymbol{\mathbf{\gamma}}_{j}\rVert_{2}$ against the number of non-zero entries in $\boldsymbol{\mathbf{\Gamma}}$ and $\{\boldsymbol{\mathbf{B}}_{h}\}$ The shape of the plot verifies the theoretical results of Section 4, that the $\ell_{2}$ error scales with the square-root of the sparsity levels.
@@
6 GBM eQTL Reanalysis
Glioblastoma multiforme (GBM) is the most malignant type of brain cancer and patient prognosis is typically very poor. Although there has been research on the genetic signaling pathways involved in the proliferation of GBM, it remains largely incurable; see Hanif et al. [2017] for a survey. It is important to understand the conditional independence structure of genes involved in GBM in order to discover new drug therapies [Kwiatkowska et al., 2013]. Our estimated graphs describe the conditional independence of co-expressions in a gene network; hence, we refer to estimated networks and effects of SNPs on this network.
We reanalyze a GBM eQTL data set that was reported in Zhang and Li [2022]. The data set contains microarray and SNP profiling data of $n=401$ GBM patients from the REMBRANDT trial (GSE108476) We use the expression levels of $p=73$ genes known to belong to the human glioma pathway according to the Kyoto Encyclopedia of Genes and Genomes (KEGG) database [Kanehisa and Goto, 2000]; the genes and pathways are detailed in Table S3. We also consider $q=118$ SNPs that are local to these $73$ genes. The SNPs are binary-coded, with 00 indicating homozygous major alleles at that locus and $1$ otherwise. Our data set is slightly different from the one used in Zhang and Li [2022]; we have a larger cohort size (compared to $n=178$ in Zhang and Li 2022) and do not include age and sex as covariates.
We ran cspine on the data set using $5$-fold cross-validation over $\alpha_{s}=0.1,0.2,\dotsc,1$, and $100$ values of $\lambda_{0}$. Our method identified $56$ SNPs that potentially modify expressions in the network. However, many of the identified edges in these networks have small weights. Since cross-validation tends to select dense models, we focus on interpreting those entries of $\{\tilde{\boldsymbol{\mathbf{B}}}_{h}\}$ above $0.005$ in magnitude. With this threshold, $10$ SNPs are estimated to have effects on the network; see Table S4.
We also ran RegGMM on the data set for comparison, using $5$-fold cross-validation for a path of $100$ $\lambda$ parameters and $10$ mixture parameters. RegGMM identified $16$ SNPs with nonzero effects on the network. However, similarly to cspine, many of the estimated edges have very small weights. We choose to threshold these edges as well.
Figure 3 shows the estimated population (covariate-independent) network from cspine and RegGMM. It can be seen that many estimated edges overlap and that cspine estimates a somewhat denser network. For example, SHC4 and CALML4 are highly connected nodes in both population networks.

### v2

@@ 5.1 Data generation
For each simulation setting, we generate a sparse coefficient matrix $\boldsymbol{\mathbf{\Gamma}}\in\mathbb{R}^{p\times q}$ by independently setting each entry $\Gamma_{jk}$ to be nonzero with probability $0.3$, with nonzero entries drawn from $N(0,1)$. The matrix $\boldsymbol{\mathbf{\Gamma}}$ is then rescaled so that the signal-to-noise ratio is approximately $1$ for each observation.
We generate the baseline precision matrix $\boldsymbol{\mathbf{B}}_{0}$ using a preferential attachment algorithm [barabasi] with power parameter equal to $1$. To introduce covariate-dependent effects, we select $q_{e}=5$ covariates to have nonzero influence and construct the matrices $\boldsymbol{\mathbf{B}}_{h}$ for $h\in[q_{e}]$ as Erdős–R’enyi graphs [erdos] with edge probability $v_{e}=0.01$. This choice ensures that the covariate-dependent graphs are sparser than the population graph [clauset].
Given the graph structure, the nonzero entries of $\boldsymbol{\mathbf{B}}_{h}$} are generated independently from the distribution $\mathrm{Unif}([-0.5,-0.35]\cup[0.35,0.5])$. To ensure diagonal dominance of the resulting precision matrix, each row $j$ is scaled by $\sum_{h}\sum_{k\neq j}\lvert\beta_{jkh}\rvert\times 1.5$ followed by symmetrization of each $\boldsymbol{\mathbf{B}}_{h}$ via averaging $\beta_{jkh}$ and $\beta_{kjh}$. This construction mirrors the data-generating mechanism in zhangli.
Among the $q$ covariates, half are generated as binary variables and half are drawn independently from $\mathrm{Unif}(0,1)$. The continuous covariates are then standardized across the $n$ observations to have zero mean and unit variance. For each observation $i\in[n]$, we construct matrix $\boldsymbol{\mathbf{B}}_{0}+\sum_{h=1}^{q}\boldsymbol{\mathbf{B}}_{h}u_{h}^{(i)}$.
5.2 Model specification and experimental setup
We evaluate both methods under two data-generating specifications. Recall that zhangli adopt the model (3), while our method is based on the alternative parametrization (9). In both models, we enforce all entries of $\operatorname{diag}(\boldsymbol{\mathbf{\Omega}}^{(i)})$ to be $1$, so that $\sigma^{2}_{\varepsilon_{j}}=1$. Under (3), we set $\boldsymbol{\mathbf{\mu}}^{(i)}=\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}}^{(i)}$. Under our specification (8) and (9), referred to as “natural model”, we set $\boldsymbol{\mathbf{\mu}}^{(i)}=[\boldsymbol{\mathbf{\Omega}}^{(i)}]^{-1}\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}}^{(i)}$. For both models, we generate responses as $\boldsymbol{\mathbf{x}}^{(i)}\sim N_{p}(\boldsymbol{\mathbf{\mu}}^{(i)},[\boldsymbol{\mathbf{\Omega}}^{(i)}]^{-1})$ for $i\in[n]$.
We consider two regimes: $(p,q)=(25,50)$ and $(25,100)$. For each setting, we generate $100$ independent datasets with sample sizes $n=200$ and $n=400$. Both RegGMM and cspine are applied to each dataset. For a fair comparison, both methods employ two tuning parameters selected via $5$-fold cross-validation. Specifically, using (16), we evaluate a grid consisting of $100$ values of $\lambda_{0}$ and $10$ values of the mixing parameter $\alpha_{s}\in{0.1,0.2,\dotsc,1}$. We report the mean and standard error (over $100$ independent datasets) of the following performance metrics:
  • TPR: true positive rate of detected edges across all $\boldsymbol{\mathbf{B}}_{h}$};
  • $\text{TPR}_{\mathrm{pop}}$: true positive rate for the population graph $\boldsymbol{\mathbf{B}}_{0}$;
  • $\text{FPR}_{\mathrm{pop}}$: false positive rate for $\boldsymbol{\mathbf{B}}_{0}$;
  • $\text{TPR}_{\mathrm{cov}}$: true positive rate for the covariate-dependent graphs $\boldsymbol{\mathbf{B}}_{h}$, $h\in[q]$;
  • $\boldsymbol{\mathbf{\beta}}_{\mathrm{err}}=\sum_{j=1}^{p}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{j}-\boldsymbol{\mathbf{\beta}}_{j}\rVert_{2}$: nodewise estimation error;
Additional metrics are reported in Section S2. In all simulation settings considered, the overall false positive rate of detected edges is below $0.005$ and is therefore not reported.
5.3 Results
@@
As expected, Table 1 shows that cspine outperforms RegGMM under the natural model across all metrics, where cspine is correctly specified while RegGMM is misspecified. Table 2 presents results under the setting (3), in which RegGMM is correctly specified and cspine is misspecified. Despite this misspecification, cspine remains competitive with RegGMM in both edge detection and estimation error, due to the convex formulation of the nodewise regressions.
$n$ $(p,q)$ Method TPR $\text{TPR}_{\text{pop}}$ $\text{FPR}_{\text{pop}}$ $\text{TPR}_{\text{cov}}$ $\boldsymbol{\mathbf{\beta}}_{\text{err}}$
200 (25, 50) cspine $\mathbf{0.73}$ $(0.08)$ $\mathbf{0.89}$ $(0.07)$ $\mathbf{0.07}$ $(0.02)$ $\mathbf{0.45}$ $(0.21)$ $\mathbf{5.66}$ $(0.50)$
RegGMM $0.62$ $(0.09)$ $0.80$ $(0.10)$ $0.16$ $(0.06)$ $0.30$ $(0.18)$ $7.43$ $(0.61)$
(25, 100) cspine $\mathbf{0.66}$ $(0.10)$ $\mathbf{0.82}$ $(0.09)$ $\mathbf{0.05}$ $(0.02)$ $\mathbf{0.40}$ $(0.24)$ $\mathbf{5.98}$ $(0.55)$
RegGMM $0.55$ $(0.08)$ $0.71$ $(0.08)$ $0.14$ $(0.05)$ $0.28$ $(0.19)$ $7.73$ $(0.56)$
400 (25, 50) cspine $\mathbf{0.77}$ $(0.10)$ $\mathbf{0.96}$ $(0.04)$ $\mathbf{0.09}$ $(0.02)$ $\mathbf{0.52}$ $(0.24)$ $\mathbf{4.50}$ $(0.59)$
RegGMM $0.67$ $(0.08)$ $0.86$ $(0.07)$ $0.17$ $(0.07)$ $0.41$ $(0.20)$ $5.99$ $(0.55)$
(25, 100) cspine $\mathbf{0.72}$ $(0.09)$ $\mathbf{0.94}$ $(0.05)$ $\mathbf{0.08}$ $(0.02)$ $\mathbf{0.44}$ $(0.21)$ $\mathbf{4.96}$ $(0.60)$
RegGMM $0.62$ $(0.08)$ $0.85$ $(0.08)$ $0.22$ $(0.07)$ $0.32$ $(0.17)$ $6.80$ $(0.58)$
@@
Table 1: Mean and standard error of performance metrics over $100$ data sets generated under specification (9), where cspine is correctly specified and RegGMM is misspecified.
$n$ $(p,q)$ Method TPR $\text{TPR}_{\text{pop}}$ $\text{FPR}_{\text{pop}}$ $\text{TPR}_{\text{cov}}$ $\boldsymbol{\mathbf{\beta}}_{\text{err}}$
200 (25, 50) cspine $\mathbf{0.53}$ $(0.06)$ $\mathbf{0.77}$ $(0.08)$ $\mathbf{0.12}$ $(0.02)$ $0.09$ $(0.10)$ $\mathbf{7.70}$ $(0.29)$
RegGMM $0.50$ $(0.07)$ $0.70$ $(0.10)$ $0.14$ $(0.05)$ $\mathbf{0.12}$ $(0.11)$ $8.02$ $(0.41)$
(25, 100) cspine $\mathbf{0.44}$ $(0.07)$ $\mathbf{0.67}$ $(0.09)$ $\mathbf{0.09}$ $(0.02)$ $0.04$ $(0.06)$ $\mathbf{7.97}$ $(0.32)$
RegGMM $0.39$ $(0.07)$ $0.59$ $(0.12)$ $0.19$ $(0.07)$ $\mathbf{0.05}$ $(0.06)$ $8.80$ $(0.45)$
400 (25, 50) cspine $\mathbf{0.60}$ $(0.05)$ $\mathbf{0.92}$ $(0.05)$ $0.18$ $(0.03)$ $0.12$ $(0.10)$ $6.71$ $(0.28)$
RegGMM $0.59$ $(0.07)$ $0.80$ $(0.10)$ $\mathbf{0.14}$ $(0.05)$ $\mathbf{0.28}$ $(0.15)$ $\mathbf{6.64}$ $(0.42)$
(25, 100) cspine $\mathbf{0.55}$ $(0.06)$ $\mathbf{0.88}$ $(0.07)$ $\mathbf{0.15}$ $(0.02)$ $0.12$ $(0.10)$ $\mathbf{6.83}$ $(0.26)$
RegGMM $\mathbf{0.55}$ $(0.07)$ $0.81$ $(0.09)$ $0.22$ $(0.07)$ $\mathbf{0.19}$ $(0.14)$ $7.15$ $(0.42)$
@@
Table 2: Performance metrics over $100$ data sets under specification (3), where RegGMM is correctly specified and cspine is misspecified.
We next provide numerical verification of the error bound in Theorem 1 through simulation. We fix $p=25$, $q=10$, and $n=300$. In the first set of experiments, we fix a collection $\{\boldsymbol{\mathbf{B}}_{h}\}$ and generate $300$ matrices $\boldsymbol{\mathbf{\Gamma}}$ of varying sparsity levels, using each to generate one data set. In the second set of experiments, we fix a sparse $\boldsymbol{\mathbf{\Gamma}}$ and generate data sets from $300$ different collections of $\{\boldsymbol{\mathbf{B}}_{h}\}$ with varying sparsity levels. In both settings, the data are generated according to specification (9).
In Figure 2, we plot the total estimation error $\boldsymbol{\mathbf{\gamma}}_{\text{err}}+\boldsymbol{\mathbf{\beta}}_{\text{err}}$, where $\boldsymbol{\mathbf{\gamma}}_{\text{err}}=\sum_{j=1}^{p}\lVert\hat{\boldsymbol{\mathbf{\gamma}}}_{j}-\boldsymbol{\mathbf{\gamma}}_{j}\rVert_{2}$, against the sparsity level of $s_{\boldsymbol{\mathbf{\Gamma}}}$ on the left panel, and against the sparsity level of $\{\boldsymbol{\mathbf{B}}_{h}\}$ on the right panel. The observed trend is consistent with the theoretical estimation error in Theorem 1, showing that the $\ell_{2}$ error scales with the square root of the sparsity level.
Finally, we illustrate the advantage of jointly optimizing over $\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\beta}}$, in contrast to the two-stage method of RegGMM. In Figure 3, we plot $\boldsymbol{\mathbf{\beta}}_{\text{err}}$ against $s_{\boldsymbol{\mathbf{\Gamma}}}$ for both methods under model specification (3) of zhangli. The data are generated with $p=25$, $q=50$, and $n=300$, with $300$ replications across varying levels of $s_{\boldsymbol{\mathbf{\Gamma}}}$.
We observe that the estimation error of RegGMM increases as $\boldsymbol{\mathbf{\Gamma}}$ becomes denser. This is consistent with the theoretical finding in zhangli, where accurate estimation of $\{\boldsymbol{\mathbf{B}}_{h}\}$ in the second stage depends on controlling the first-stage estimation error $\lVert\hat{\boldsymbol{\mathbf{\Gamma}}}-\boldsymbol{\mathbf{\Gamma}}\rVert_{F}$, which becomes more challenging as $\boldsymbol{\mathbf{\Gamma}}$ becomes less sparse. In constrast, cspine exhibits stable performance across different sparsity levels of $\boldsymbol{\mathbf{\Gamma}}$, highlighting the advantage of the joint convex formulation in (14), which avoids error propagation from a separate centering stage.
@@
Figure 3: Estimation error of cspine and RegGMM over $300$ data sets plotted against the number of non-zero entries in $\boldsymbol{\mathbf{\Gamma}}$ with a small horizontal jitter applied. cspine is seen to have better scaling properties in the sparsity of $\boldsymbol{\mathbf{\Gamma}}$. Data were generated via the model specification (3) of zhangli.
6 GBM eQTL analysis
Glioblastoma multiforme (GBM) is the most malignant type of brain cancer and patient prognosis is typically very poor. Although there has been research on the genetic signaling pathways involved in the proliferation of GBM, it remains largely incurable; see glio-hanif for a survey. It is important to understand the conditional independence structure of genes involved in GBM in order to discover new drug therapies [glio-kwiakowska]. Our estimated graphs describe the conditional independence of co-expressions in a gene network; hence, we refer to estimated networks and effects of SNPs on this network.
We reanalyze a GBM eQTL data set that was reported in zhangli. The data set contains microarray and SNP profiling data of $n=401$ GBM patients from the REMBRANDT trial (GSE108476). We use the expression levels of $p=73$ genes known to belong to the human glioma pathway according to the Kyoto Encyclopedia of Genes and Genomes (KEGG) database [KEGG]; the genes and pathways are detailed in Table S3. We also consider $q=118$ SNPs that are local to these $73$ genes. The SNPs are binary-coded, with 0 indicating homozygous major alleles at that locus and $1$ otherwise. Our data set is slightly different from the one used in zhangli; we have a larger cohort size (compared to $n=178$ in zhangli) and do not include age and sex as covariates.
We ran cspine on the data set using $5$-fold cross-validation over $\alpha_{s}=0.1,0.2,\dotsc,1$, and $100$ values of $\lambda_{0}$ in (16). Since cross-validation tends to select relatively dense models, we focus on entries of $\tilde{\boldsymbol{\mathbf{B}}}_{h}$} with magnitude exceeding $0.01$. Under this threshold, our method identifies $28$ SNPs that potentially modify expression patterns in the network. For comparison, we also apply RegGMM to the dataset, using $5$-fold cross-validation over a grid of $100$ $\lambda_{0}$ values and $10$ values of $\alpha_{s}$. After applying the same thresholding procedure, RegGMM identifies $12$ SNPs with nonzero effects on the network. The results of both methods are summarized in Table S4.
Figure 4 displays the estimated population networks, corresponding to $\boldsymbol{\mathbf{B}}_{0}$, obtained from cspine (left) and RegGMM (right). The two networks share many common structures; for example, E2F3, an oncogenic transcription factor, and CALML4, a protein-coding gene, are highly connected nodes in both population networks. RegGMM produces a substantially denser population network, with $750$ total edges compared to $476$ edges identified by cspine.